### v1

# Algebraic Control: Complete Stable Inversion with Necessary and Sufficient Conditions

Abstract
Recent advances in learning-based control have increased interest in stable inversion to meet growing performance demands. Here, we establish necessary and sufficient conditions for stable inversion, addressing challenges in non-minimum phase, non-square, and singular systems. An $\mathcal{H}_{\infty}$-based algebraic approximation is introduced for near-perfect tracking without preview. Additionally, we propose a novel robust control strategy combining the nominal model with dual feedforward control to form a feedback structure. Numerical comparison demonstrates the approach’s effectiveness.
Index Terms:
Stable inversion, non-minimum phase systems, robustness, algebraic control, multivariable systems
I Introduction
Recent advancements in trajectory planning, autonomous vehicles, and robotics have increased interest in stable inversion to meet performance demands [1]. Learning-based control methods have further broadened research into inversion [2, 3], emphasizing the need to re-examine the conditions for stable inversion.
The study of stable inversion began with Brockett’s introduction of functional reproducibility in 1965 [4]. Silverman extended these concepts to multivariable systems in 1969 [5], followed by geometric formulations introduced by Basile and Marro [6] and applications in reduced-order control by Moylan [7]. The stable inversion in discrete-time systems was first addressed by Tomizuka [8], while Hunt explored non-causal inversion techniques [9].
Recent advancements include stable inversion for SISO affine systems [10], MIMO square systems in continuous time [11] and discrete time [12], and approximate stable inversion [13]. Specialized approaches, such as strong inversion for minimum-phase systems [14] and geometric methods for convolution-based inversion [15], have further expanded the field.
As the field evolves, the need for system classification and tailored solutions becomes evident. The primary objective of inversion-based approaches is to determine a bounded input that accurately reproduces the desired or given output. However, this goal poses challenges, particularly with non-minimum phase systems, which complicate the establishment of stable inputs under causality without infinite pre-actuation [9]. Furthermore, stable inversion can impact tracking performance and stability in the presence of uncertainties [16]. Merging learning-based control with stable inversion addresses these issues [10, 17].
Stable inversion methods have been applied across various domains, including iterative learning control [18], trajectory tracking for autonomous systems [19], and optimal channel equalization [20].
In this paper, we extend the concept of Zero Phase Error Tracking (ZPET) [8] to continuous-time systems. Our study focuses on determining the necessary and sufficient conditions for stable inversion in non-square, non-minimum phase, and uncertain systems. The main contributions are as follows:
  • Revisiting the stable inversion problem using a similar algebraic setting in Exact Model Matching [21] yet explicitly stating the output structures for necessary and sufficient conditions, which also hold for non-minimum phase systems. Compared to [14, 13, 11], the exact inversion conditions for non-square and even singular systems are also established.
  • Proposing an algebraic approximation method based on $\mathcal{H}_{\infty}$ theory when perfect tracking conditions are not met. This method allows for near-perfect tracking without the need for preview or pre-actuation approaches, which differentiates it from the methods used in [9, 15].
  • Ensuring robustness guarantees without the need for learning mechanisms, complementing the strategy in [10]. Furthermore, we establish upper bounds on achievable performance for non-minimum phase uncertain systems, extending the results presented in [13].
Our ultimate goal is to establish the necessary and sufficient conditions for stable inversion for non-minimum and uncertain systems in a causal manner, providing a unified framework to address the challenges of stable inversion.
The remainder of the paper is organized as follows: Section II provides the algebraic preliminaries and problem formulation. Section III presents the main results, including exact and almost necessary and sufficient conditions, as well as robustness. Section IV illustrates numerical examples to validate the theory. Section V concludes the paper.
II Preliminaries
II-A Algebraic Preliminaries
Let $\mathbb{R}$ be a field of real numbers. Consider the set of all polynomials in $s$ with coefficients in $\mathbb{R}$. This set forms a commutative ring over $\mathbb{R}$ and is denoted by $\mathbb{R}[s]$. If this ring, say $\mathcal{R}$, has an identity element and no zero divisors, then $\mathcal{R}$ is an integral domain. Note that $\mathbb{R}[s]$ is also an Euclidian Domain. Moreover, the set of all such rational functions in $s$ over $\mathbb{R}$ forms a field, denoted by $\mathbb{R}(s)$ [22], and $\mathbb{R}[s]\subset\mathbb{R}(s)$ holds. The sets of $n_{y}\times n_{u}$ matrices with elements in $\mathbb{R}$, $\mathbb{R}[s]$, and $\mathbb{R}(s)$ are denoted by $\mathbb{R}^{n_{y}\times n_{u}}$, $\mathbb{R}[s]^{n_{y}\times n_{u}}$, and $\mathbb{R}(s)^{n_{y}\times n_{u}}$ respectively. Any $n_{y}\times n_{u}$ matrix, say $\grave{P}(s)\in\mathbb{R}[s]^{n_{y}\times n_{u}}$, over $\mathcal{R}$ can be factorized (see [23, Theorem 2.1]) as
$\grave{P}(s)$ $={\color[rgb]{0,0,0}U_{R}(s)S_{m}(s)V_{R}(s)}$
$=[{U}_{R1}\ {U}_{R2}]\begin{bmatrix}\Lambda(s)&\textbf{0}_{r\times(n_{u}-r)}\\ \textbf{0}_{(n_{y}-r)\times r}&\textbf{0}_{(n_{y}-r)\times(n_{u}-r)}\end{bmatrix}{\begin{bmatrix}{V}_{R1}\\ {V}_{R2}\end{bmatrix}}$ (1)
where $U_{R}(s)\in\mathbb{R}[s]^{n_{y}\times n_{y}}$ and $V_{R}(s)\in\mathbb{R}[s]^{n_{u}\times n_{u}}$ are $\mathcal{R}$-Unimodular matrices, $\Lambda(s)=\text{diag}[d_{1}(s)\ \dots\ d_{r}(s)]$ with unique monic $d_{i}(s)\in\mathcal{R}$ such that $d_{i}$ divides $d_{i+1}$. The matrix $S_{m}(s)\in\mathbb{R}[s]^{n_{y}\times n_{u}}$ is called as Smith Normal form of $\grave{P}$.
Let $V$ be a vector space over the field $\mathbb{R}(s)$ with dimension $k$, consisting of $n$-tuples such that a basis of vector polynomials can always be found for $V$ [page 22, [24]]. A minimal basis of $V$ is defined as a $k\times n$ polynomial matrix $P_{m}$ [22]. Adapted Forney’s Theorem [22, Section 3(4.)] states that if $y=xP_{m}$ is a polynomial $n$-tuple, then $x$ must be a polynomial $k$-tuple.
For the rank notation, akin to [25], consider $P$ belongs to $\mathbb{R}(s)^{n_{y}\times n_{u}}$. Then, the rank of $P$ is defined as the maximum size of any linearly independent subset of its columns in the field $\mathbb{R}(s)$, denoted by $\text{rank}_{\mathbb{R}(s)}(P)$, where $\text{rank}_{\mathbb{R}(s)}(P)\neq\text{rank}_{\mathbb{R}}(P)$.
Suppose $\text{rank}_{\mathbb{R}(s)}(P)=r$, where $1\leq r\leq\min(n_{u},n_{y})$. Define $J=\{j_{1},\dots,j_{r}\}\subseteq\{1,\dots,n\}$ as an ordered indexed set corresponding to $P(s)$’s linearly independent columns. We then define the matrix $L$ as
${L}=[p_{j_{i}}\ \dots\ p_{j_{r}}]\in\mathbb{R}(s)^{n_{y}\times r}\,$ (2)
where $p_{j_{k}}$ denotes the $j_{k}$-th column of $P$. By construction, $L$ has full column rank $r$, forming a basis for the column space of $P$ while preserving its span with a minimal set of linearly independent columns. The image of $P$ over $\mathbb{R}(s)$ is then:
$\text{Im}_{\mathbb{R}(s)}(P)=\left\{\sum_{i=1}^{r}c_{i}p_{i}:c_{i}\in\mathbb{R}(s),p_{i}\in{L}\right\}\subseteq\mathbb{R}(s)^{n_{y}}.$ (3)
Some further notations throughout the paper are: (.$\|(.)(t)\|_{\infty}\triangleq\text{ess sup}_{t}|(.)(t)|$, any complex number can be expressed as (.)±$\Re(.)\pm j\Im(.)\in\mathbb{C}$, $\sigma$ represents singular values. Time domain square-integrable functions are denoted by $\mathcal{L}_{2}(-\infty,\infty)$. Its causal subset is given by $\mathcal{L}_{2}[0,\infty)$. For the frequency domain, including at $\infty$, $\mathcal{L}_{2}(j\mathbb{R})$ represents square-integrable functions on $j\mathbb{R}$, and $\mathcal{L}_{\infty}(j\mathbb{R})$ denotes bounded functions on $\Re(s)=0$. All these functions in Lebesgue spaces may be either matrix-valued or scalar. Then, $\mathcal{R}\mathcal{H}_{\infty}$ denotes the set of real rational $\mathcal{L}_{\infty}(j\mathbb{R})$ functions analytic in $\Re(s)>0$, $(.)^{\dagger}$ represents pseudo-inverse yielding unit matrix under multiplication, $(.)^{*}$ denotes complex conjugates transpose, and .,$\langle$.,$\rangle$ . , . ⟩ denotes the inner product. Now, among various definitions for multivariable zeros of a Transfer Function Matrix (TFM), $P(s)$, we adopt the definition given in [26, chapter 4.5.3] as
$\text{Z}_{P}$$\triangleq\{z\in\mathbb{C}:P(z)u_{z}=0y_{z}\}$ (4)
$\text{Z}_{R}$$\triangleq\{z\in\text{Z}_{P}:\Re(z)>0\}\subseteq\text{Z}_{P}$ (5)
where $\text{Z}_{R}$ defines the RHP zeros making the system non-minimum phase, $u_{z},y_{z}$ (can be obtained via singular value decomposition (SVD) of $P(z)=U\Sigma V^{*}$) are normalized input and output zero directions respectively.
II-B Problem Statement
Consider the general representation of a MIMO linear time-invariant (LTI) system, $\Xi:\mathbb{U}\times\mathbb{X}\to\mathbb{Y}$, as
$\Xi(u,x_{0}):\begin{cases}\dot{x}(t)&=Ax(t)+Bu(t)\\ y(t)&=Cx(t),\ x_{0}\triangleq x(0)\neq 0\end{cases}$ (6)
where $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times n_{u}}$, $C\in\mathbb{R}^{n_{y}\times n}$, $x,x_{0}\in\mathbb{X}\subseteq\mathbb{R}^{n},x(t)\in\mathcal{L}_{2}[0,\infty),y\in\mathbb{Y}\subset\mathbb{R}^{n_{y}},y(t)\in\mathcal{L}_{2}[0,\infty),u\in\mathbb{U}\subset\mathbb{R}^{n_{u}}$, and $u(t)\in\mathcal{L}_{2}[0,\infty)$. Moreover, the TFM representation of $\Xi({\color[rgb]{0,0,0}u},0)\triangleq C(sI-A)^{-1}B=P(s)$ and since all real rational strictly proper transfer matrices with no poles on the imaginary axis form $\mathcal{RL}_{2}(j\mathbb{R})\subset\mathcal{L}_{2}(j\mathbb{R})$ [27, page 48], $P(s)$ belongs to $\mathcal{L}_{2}(j\mathbb{R})\cap\mathbb{R}(s)^{n_{y}\times n_{u}}$. The system, $\Xi$, is; P1) a non-minimum phase s.t. $Z_{R}\neq\emptyset$, P2) either $n_{y}=n_{u}$ (square) or $n_{y}\neq n_{u}$ (nonsquare), and casual ($t>0$).
Assumption 1.
The system given in (6) is minimal.
Problem 1.
Under Assumption 1, let $y_{1}(t)$ be an observed output trajectory of the system $\Xi$ in (6) resulting from some unknown initial condition $\hat{x}_{0}$ and control input $\hat{u}$.
Now, given trajectory $y_{1}(t)$ and initial condition $x_{0}$, the left inverse, $\Xi_{i}^{\ell}:\mathbb{Y}\times\mathbb{X}\to\mathbb{U}$, is defined such that
$u_{inv}(t)=\Xi_{i}^{\ell}(y_{1}(t),x_{0}),\quad\|u_{inv}(t)\|_{\infty}<\infty.$ (7)
The error is
$e(t)$ $\triangleq y_{1}(t)-\Xi(u_{inv}(t),x_{0})$ (8)
The inverse, $\Xi_{i}^{\ell}$, is classified based on $e(t)$ as follows:
  • a.
    Stable exact inverse: if $e(t)=0$ for all $t>0$.
  • b.
    Stable approximate inverse: if $\|e(t)\|_{\infty}<\infty$.
For approximate inversion, among all possible approximations, our goal is to find $\Xi_{i}^{\ell}$ that minimizes the error $e(t)$ with $\|u_{inv}(t)\|_{\infty}<\infty$ and satisfying
$\|e(t)\|\leq\alpha\|e(0)\|e^{-\beta t}+\varphi,\quad\forall t>0,$ (9)
for some $\alpha,\beta>0$, and $\varphi\geq 0$.
Remark 1.
The terminology in Problem 1, adopted from [28], allows $y_{1}(t)$ to represent desired trajectories, as in [12]. Assumption 1 and $\|e(t)\|_{\infty}<\infty$ ensure Lyapunov stability for the error dynamics in (8). The objective of minimizing the error aligns with [13], where $\varphi$ in (9) accounts for bounded yet inevitable errors, including steady-state errors due to system limitations.
To solve Problem 1.a algebraically, given $y_{1}(t)$, define $y_{ic}(t)\triangleq\Xi(0,x_{0})$, the trajectory caused by the known initial conditions. To isolate $y_{ic}(t)$ in $y_{1}(t)$, we take the Laplace transform (see Paley-Wiener Th., [29, p.104] for the existence) yielding
[…]
$\quad\quad\Rightarrow\ P(s)U(s)=Y(s)\quad s.t.$ (11)
$\begin{bmatrix}P_{11}(s)&\cdots&P_{1n_{u}}(s)\\ \vdots&\ddots&\vdots\\ P_{n_{y}1}(s)&\cdots&P_{n_{y}n_{u}}(s)\end{bmatrix}\begin{bmatrix}U_{1}(s)\\ \vdots\\ U_{n_{u}}(s)\end{bmatrix}=\begin{bmatrix}Y_{1}(s)\\ \vdots\\ Y_{n_{y}}(s)\end{bmatrix}.$ (12)
From this point on, similar to the [21, Theorem 8.5.2], Problem 1.a reduces to algebraically solving the rational matrix equation.
Remark 2.
The algebraic solution of (11) is not restricted solely to the framework of (10). It can also be expressed through describing functions [30, Sec III.C]. Furthermore, the recent work in [31] presents the Koopman resolvent, a Laplace-domain representation for specific nonlinear dynamics. This development emphasizes the broader potential for algebraic solutions in strongly nonlinear systems.
III Main Results
III-A Left Inverse
Assumption 1 and derivation of (10) show that all rational function entries of $P(s)$ and $Y(s)$ in (11) shares the same greatest common denominator, $\Delta=\det(sI-A)\in\mathbb{R}[s]$. To cancel out $1/\Delta$, we multiply (11) by $\Delta$, yielding
$\grave{P}(s)U(s)=\grave{Y}(s)$ (13)
where $\grave{P}(s)\triangleq U_{R}(s)S_{m}(s)V_{R}(s)U(s)\in\mathbb{R}[s]^{n_{y}\times n_{u}}$, $U(s)\in\mathbb{R}(s)^{n_{u}}$, and $\grave{Y}(s)\in\mathbb{R}[s]^{n_{y}}$. Then, using the invariant factor theorem given by (II-A) on $\grave{P}(s)$ in (13), we have
$U(s)$ $=[\hat{V}_{R1}\ \hat{V}_{R2}]\begin{bmatrix}\Lambda^{-1}(s)&\textbf{0}_{(n_{u}-r)\times r}\\ \textbf{0}_{(n_{u}-r)\times r}&\textbf{0}_{(n_{u}-r)\times(n_{y}-r)}\end{bmatrix}\begin{bmatrix}\hat{U}_{R1}\\ \hat{U}_{R2}\end{bmatrix}\grave{Y}(s)$
$\qquad+(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})\kappa$
$=\Biggr{(}[\hat{V}_{R1}\ \hat{V}_{R2}]\begin{bmatrix}\Lambda^{-1}(s)&\textbf{0}_{(n_{u}-r)\times r}\\ \textbf{0}_{(n_{u}-r)\times r}&\textbf{0}_{(n_{u}-r)\times(n_{y}-r)}\end{bmatrix}\begin{bmatrix}\hat{U}_{R1}\\ \hat{U}_{R2}\end{bmatrix}$
$\qquad+(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})\bar{\kappa}\Biggr{)}\grave{Y}(s)=\Xi_{i}^{\ell}(s)\grave{Y}(s)$ (14)
where $\Lambda^{-1}(s)=\text{diag}[{1}/{d_{1}(s)}\ \dots 1/{d_{r}(s)}]\in\mathbb{R}(s)^{r\times r}$, $\hat{U}_{R}(s)\in\mathbb{R}[s]^{n_{y}\times n_{y}}$ and $\hat{V}_{R}(s)\in\mathbb{R}[s]^{n_{u}\times n_{u}}$ are $\mathcal{R}$-Unimodular matrices such that ${U}_{R}(s)\hat{U}_{R}(s)=I_{n_{y}}$ and ${V}_{R}(s)\hat{V}_{R}(s)=I_{n_{u}}$ respectively, and $\kappa,\bar{\kappa}\in\mathcal{R}\mathcal{H}_{\infty}\ (\bar{\kappa}\triangleq\kappa(\grave{Y}^{T}(s)\grave{Y}(s))^{-1}\grave{Y}^{T}(s))$ are any arbitrary vectors.
III-B Algebraic Necessary and Sufficient Conditions
In this subsection, to find the solution(s) for Problem 1.a, we will show the solvability of (11) algebraically and therefore the existence and boundedness conditions of (III-A).
Theorem 1.
Consider $\Xi$ in (6) with P1-P2 and $\Xi^{\ell}_{i}$ in (III-A). Then sufficient and necessary condition for Problem 1.a. is
$\Bigl{(}e(t)=0,\ \|u(t)\|_{\infty}<\infty\Bigl{)}\iff$ $\Bigl{(}Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)\Bigl{)}\text{ and }$
$\Bigl{(}Y(z)=0,\ \forall z\in Z_{R}\Bigl{)}.$
Proof.
Note that $e(t)=0$ implies (11) and (13) hold and vice-verca. We can now proceed to the proof for both directions.
[…]
(1:) $e(t)=0$ means (III-A) holds. Then, pre-multiplying (III-A) by $\grave{P}(s)=U_{R}(s)S_{m}(s)V_{R}(s)$ and simplifying yields
$U_{R1}\hat{U}_{R1}\grave{Y}(s)=\grave{Y}(s)$ (15)
which means $\grave{Y}(s)$ is invariant under the orthogonal projection onto $\operatorname*{Im}_{\mathbb{R}(s)}(P)$. Thus, we get $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$.
(2:) Note that $(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})$ in (III-A) is pure polynomial and $\kappa\in\mathcal{R}\mathcal{H}_{\infty}$. Therefore, $(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})\kappa$ does not contribute any unstable solution(s) to $U(s)$. So, considering only the $\hat{V}_{R1}(s)\Lambda^{-1}(s)\hat{U}_{R1}(s)\grave{Y}(s)$ is enough for stability. Lets consider two cases on $\grave{P}(s)$; C1: $\operatorname*{rank}_{\mathbb{R}(s)}(\grave{P})=\min(n_{u},n_{y})$, C2: $\operatorname*{rank}_{\mathbb{R}(s)}(\grave{P})<\min(n_{u},n_{y})$. For C2, re-writing (III-A) and ignoring $(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})\kappa$, we get
$U(s)$ $=\hat{V}_{R1}\Lambda^{-1}\hat{U}_{R1}\grave{Y}(s)$
$\Lambda\hat{V}^{\ell}_{R1}U(s)$ $=\hat{U}_{R1}\grave{Y}(s)$
[…]
where there always exist $\hat{V}^{\ell}_{R1}$ satisfying $\hat{V}^{\ell}_{R1}\hat{V}_{R1}=I_{r}$. Now assume that either $\grave{P}(s)$ or $U_{R1}\Lambda\hat{V}^{\ell}_{R1}$ is not a minimal basis. The minimal basis $\grave{P}_{m}(s)$ of an $n_{y}$-dimensional vector space of $n_{u}$-tuples over $\mathbb{R}[s]$ for (13) is given by $\grave{P}_{m}(s)=L_{T}(s)\grave{P}(s)$ for condition C1, and $\grave{P}_{m}(s)=L_{T}(s)U_{R1}\Lambda\hat{V}^{\ell}_{R1}$ for condition C2. Here, the transformation $L_{T}(s)\in\mathbb{R}(s)^{n_{y}\times n_{y}}$, which yields the minimal basis, exists with $\det(L_{T})\neq 0$ (see [22, Section 2]). Then, left multiplying $L_{T}(s)$ to (13) or (16) results in
$\grave{P}_{m}(s)U(s)=L_{T}(s)\grave{Y}(s)$ (17)
where the equality and solution are preserved. Now we can re-write $L_{T}(s)$ as $L_{T}(s)=\frac{1}{\Delta_{L}}L^{\prime}_{T}$ s.t. $L^{\prime}_{T}\in\mathbb{R}[s]^{n_{y}\times n_{y}}$, and the scalar $\Delta_{L}\in\mathbb{R}[s]$. So, multiplying (17) by $\Delta_{L}$ to yields
$\grave{P}_{m}(s)\Delta_{L}U(s)=L^{\prime}_{T}(s)\grave{Y}(s)$ (18)
[…]
$\underbrace{y^{*}_{z}\grave{P}_{m}(z)}_{=0}\Delta_{L}(z)U(z)=\underbrace{y^{*}_{z}L^{\prime}_{T}(z)}_{\neq 0}\grave{Y}(z)$ (19)
Since the left-hand side of (19) becomes zero and since $\det(L_{T})\neq 0$ and $1/\Delta_{L}(s)$ is stable, the only way to satisfy the equality is $\grave{Y}(z)=0$.
[…]
For boundedness, consider (5) and rewrite (13) as
$\grave{P}(s)U(s)=\grave{P}(s)\frac{U_{N}(s)}{U_{D}(s)}=\grave{Y}(s)$ (20)
$\grave{P}(s)U_{N}(s)=U_{D}(s)\grave{Y}(s)$ (21)
where $U_{N}(s)\in\mathbb{R}[s]^{n_{u}}$ and scalar $U_{D}(s)\in\mathbb{R}{\color[rgb]{0,0,0}[s]}$. Now define a new scalar $z^{+}(s)$ which contains all RHP zeros of $\grave{P}(s)$ as
$z^{+}(s)=\prod_{z\in Z_{R}{\color[rgb]{0,0,0}\bigcap\mathbb{R}}}(s-z)^{m_{z}}\prod_{z\in{\color[rgb]{0,0,0}Z_{R}\bigcap(\mathbb{C}\setminus\mathbb{R})}}(s-z)^{m_{z}}(s-{z}^{*})^{m_{{z}^{*}}}$ (22)
where $z^{+}(s)\in\mathbb{R}[s]\subset\mathbb{R}(s)$, $m_{z}$ and $m_{z^{\ast}}$ denote the multiplicities of $z$ and ${z}^{*}$ respectively. Thus, there exist non-zero vector $y^{\ast}_{z}\ (\text{or }u_{z})$ such that $y^{\ast}_{z}\grave{P}(z)=0$ (or $\grave{P}(z)u_{z}=0.y_{z}$) $\forall z\in Z_{R}$.
Now suppose $\det(U_{D}(\bar{z}))=0$ for some $\bar{z}$ where $\Re(\bar{z})>0$ which means $U(s)$ is unstable. Either, suppose $(s-\bar{z})\in z^{+}(s)$, then left multiplying $y^{\ast}_{\bar{z}}$ to (13) and evaluating at $s=\bar{z}$ yield
$\Biggl{(}\underbrace{(s-\bar{z})\begin{bmatrix}\ast(s)_{1}\\ \vdots\\ \ast(s)_{n_{y}}\end{bmatrix}}_{=y^{\ast}_{\bar{z}}\grave{P}(s)}\frac{U_{N}(s)}{\underbrace{(s-\bar{z})U^{\prime}_{D}(s)}_{=U_{D}(s)}}\Biggr{)}(\bar{z})\neq 0=y^{\ast}_{\bar{z}}\underbrace{\grave{Y}(\bar{z})}_{=0}$ (23)
or suppose $(s-\bar{z})\notin z^{+}(s)$ then left multiplying $y^{\ast}_{\bar{z}}$ to (21) and evaluating at $s=\bar{z}$ yield
$\underbrace{y^{*}_{\bar{z}}\grave{P}(\bar{z})}_{\neq 0}{U_{N}(\bar{z})}\neq 0={y^{*}_{\bar{z}}}\grave{Y}(\bar{z})\underbrace{U_{D}(\bar{z})}_{=0}$ (24)
(23) and (24) are leading to a contradiction. Thus $U(s)$ includes only LHP poles. The proof is now complete. ∎
Remark 3.
While the algebraic equation is inspired by Exact Model Matching for invertibility, we explicitly define necessary and sufficient conditions for the system’s output, applicable even to non-minimum phase systems by leveraging polynomial bases. Specifically, the condition $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$ defines the set of outputs reachable through the input channel, while $Y(z)=0,\ \forall z\in Z_{R}$ redefines the output on $Z_{R}$. Relaxing either condition results in losing the property “$=0$” over the $e(t)$.
Notably, for certain specific conditions, such as minimum phase systems, the findings can be applied more directly.
Corollary 1.
Assume $P(s)$ is square and full rank so that $U_{R}=U_{R1},V_{R}=V_{R1},(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})=0$, and $Z_{R}=\emptyset$, then (III-A) becomes
$U(s)={\hat{V}_{R1}\Lambda^{-1}(s)\hat{U}_{R1}}Y(s).$ (25)
Corollary 2.
Assume $P(s)$ is non-square and full rank, so that $V_{R}=V_{R1},(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})=0$, and $Z_{R}=\emptyset$, then (III-A) becomes
$U(s)$ $=\begin{bmatrix}\hat{V}_{R1}\Lambda^{-1}(s)&\textbf{0}_{n_{u}\times(n_{y}-n_{u})}\end{bmatrix}\begin{bmatrix}\hat{U}_{R1}\\ \hat{U}_{R2}\end{bmatrix}Y(s)$ (26)
$=(P^{*}(s)P(s))^{-1}P^{*}(s)Y(s)=P^{\dagger}(s)Y(s).$ (27)
III-C Almost Necessary and Sufficient Conditions
In this section, we propose approximate remedies for non-minimum phase systems. Our goal is to relax, at the cost of losing exactness, the and condition in Theorem 1, $\bigl{(}Y(z)=0,\ \forall z\in Z_{R}\bigl{)}$. This can be achieved using the $\mathcal{H}_{\infty}$-Synthesis framework.
We define a virtual loop in the sense of classical feedback structure shown by Fig.1.(a). Given the system $P(s)$ as in (11), suppose we have a virtual controller, say $K(s)\in\mathbb{R}(s)^{n_{u}\times n_{y}}$, be defined such that $u_{\upsilon}(t)=K(s)(r_{\upsilon}(t)-y_{\upsilon}(t))$. The key transfer functions within the virtual feedback loop are defined as follows: the input complementary sensitivity function, $T_{i}(s)\triangleq(I_{n_{u}}+K(s)P(s))^{-1}K(s)P(s)$; the input sensitivity function, $S_{i}(s)\triangleq(I_{n_{u}}+K(s)P(s))^{-1}$; and the loop transfer function, $L_{i}(s)\triangleq K(s)P(s)$. By breaking the loop at the output gives the output functions, $T_{o},S_{o},L_{o}$.
Theorem 2 (Internal Stability, [27]).
Under Assumption 1, the closed-loop virtual system is internally stable iff
$\begin{bmatrix}I&K\\ -P&I\end{bmatrix}^{-1}\in\mathcal{R}\mathcal{H}_{\infty}$ (28)
Lemma 1 (Youla’s Statements[26]).
Assume (28) holds. Then the following statements are true:
  1. 1.
    If $P(s)$ has a RHP zero at $z$, then $PK$,$KP$, $PK(I+PK)^{-1}$, $(I+PK)^{-1}P$, and $KP(I+KP)^{-1}$ each have a RHP-zero at $z$.
  2. 2.
    If $P(s)$ has a RHP-pole at $p$, then $PK$ and $KP$ also have a RHP-pole at $p$, while $(I+PK)^{-1}$, $K(I+PK)^{-1}$,$(I+KP)^{-1}$ and $(I+KP)^{-1}K$ each have a RHP-zero at $p$.
Now consider augmenting the performance weight ${W}_{P}$, ${W}_{P}~{}\triangleq\text{diag}[{W_{P1}}(s)\ {W_{P2}}(s)\ \dots\ {W_{Pn_{y}}}(s)]$ to form the linear fractional transformation (LFT) as
$\begin{bmatrix}z(t)\\ e_{\upsilon}(t)\end{bmatrix}={\begin{bmatrix}W_{P}&W_{P}P\\ -I&-P\end{bmatrix}}\begin{bmatrix}w(t)\\ u_{\upsilon}(t)\end{bmatrix}{\color[rgb]{0,0,0}=G\begin{bmatrix}w(t)\\ u_{\upsilon}(t)\end{bmatrix}}$ (29)
where each scalar in $W_{P}$ is as defined in [33, Equation 22], $e_{\upsilon}(t)\triangleq r_{\upsilon}(t)-y_{\upsilon}(t)$, and $z(t)\triangleq W_{P}e(t)$. The LFT is then given by $\mathcal{F}_{l}(G,K)\triangleq{G}_{11}+{G}_{12}{K}({I}-{G}_{22}{K})^{-1}{G}_{21}$ where the $\mathcal{H}_{\infty}$ control problem involves finding $K$ such that
$\operatorname*{minimize}_{K}$ $\max_{\omega}\bar{\sigma}\Bigl{(}{G}_{11}+{G}_{12}{K}({I}-{G}_{22}{K})^{-1}{G}_{21}\Bigl{)}(j\omega)$
$\operatorname*{subject\ to}2$ $\begin{bmatrix}I&K\\ -P&I\end{bmatrix}^{-1}\in\mathcal{R}\mathcal{H}_{\infty}.$ (30)
By adhering to analytical constraints (e.g., the waterbed effect, logarithmic integrals, and interpolation bounds) defined in [34] and following the procedure in [35], we ensure $\|\mathcal{F}_{l}(G,K)\|_{\infty}=\gamma<\infty$.
By letting $\epsilon_{i}\to 0$, we replace all approximate integrators in $K$ with pure integrators [36] that imposes: If $\bar{\sigma}(S_{i}(j(0,\bar{\omega}]))<\mathbb{E}[N^{2}]$, then $\bar{\sigma}(S_{i}(j(0,\bar{\omega}]))=0$ where $\mathbb{E}[N^{2}]$ denotes the expected value mean square of random noise in dB .
Theorem 3.
Consider $P(s)$ as defined in (11)-(12) with $\text{rank}_{\mathbb{R}(s)}(P)=r=n_{u}=n_{y}$, and $Z_{R}\neq\emptyset$. Define the approximate inverse $\Xi_{a}^{\ell}:\mathbb{Y}\times\mathbb{X}\to\mathbb{U}$ as
$\Xi_{a}^{\ell}(s)=S_{i}(s)K(s),$ (31)
then, by noting (10) and rewriting (7), the control input is
$u_{inv}(t)=\mathscr{L}^{-1}\{S_{i}(s)K(s)Y(s)\},$ (32)
which satisfies $\|u_{inv}(t)\|_{\infty}<\infty$ and
$\|e(t)\|\leq\alpha\|e(0)\|e^{-\beta t},\quad\forall t>0,$
for some $\alpha,\beta>0$.
Proof.
As in Corollary 1, since $(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})=0$, we have a unique solution $P^{-1}$ that yields $e(t)=0,\forall t>0$ but $\text{lim}_{t\to\infty}u_{inv}(t)=\infty$. Therefore to obtain a stable control input, we begin by using (10) and taking the Laplace transform:
$E(s)=Y(s)-P(s)U_{inv}(s).$ (33)
Next, substitute $U_{inv}(s)=S_{i}(s)K(s)Y(s)$ in (33), noting that $S_{i}(s)=(I+K(s)P(s))^{-1}$:
$E(s)=Y(s)-P(s)(I+K(s)P(s))^{-1}K(s)Y(s).$
Rewriting,
$E(s)$ $=Y(s)[I-\underbrace{(I+P(s)K(s))^{-1}P(s)K(s)}_{\triangleq T_{o}(s)}]$ (34)
$E(s)$ $=Y(s)(I-T_{o}(s))=Y(s)S_{o}(s)$ (35)
If $Y(s)$ has any unstable pole $p$, the construction of $K(s)$ via (III-C) satisfies Theorem 2 and, therefore, Lemma 1 which ensure $S_{o}(p)=0$ and $S_{i}K(p)=0$. Thus, because $S_{o}(s)\in\mathcal{R}\mathcal{H}_{\infty}$ and $S_{o}(p)=0$, it follows that $E(s)=Y(s)S_{o}(s)$ is stable and bounded (i.e., $E(s)\in\mathcal{R}\mathcal{H}_{\infty}$ ). Therefore $e(t)=\mathscr{L}^{-1}\{E(s)\}$ decays exponentially, implying $\|e(t)\|\leq\alpha\|e(0)\|e^{-\beta t},\forall t>0.$
Similarly, since $S_{i}(s)K(s)\in\mathcal{R}\mathcal{H}_{\infty}$ and $S_{i}K(p)=0$, it follows that
$u_{inv}(t)=\mathscr{L}^{-1}\{S_{i}(s)K(s)Y(s)\}$
is also bounded. This completes the proof. ∎
Corollary 3.
Control input $u_{inv}(t)$ in Theorem 3 applies to the case where $P(s)$ has more inputs than outputs (i.e., $n_{u}>n_{y}$) but full rank.
Now let us define $P_{p}(s)\in\mathbb{R}(s)$ which is a scalar approximation of MIMO $P(s)$ in (11). To do it first consider one entry of $P(s)$ for $n_{u}<n_{y}$ (taking an average of an entire column is also an option). Because of Assumption 1, the characteristic equation of $P_{p}(s)$, $\Delta_{p}$, equals $\Delta$. Then, if for any RHP zeros of $P(s)$ does not a RHP zeros of $P_{p}(s)$, zero augmentation as $P_{p}(s)=z^{+}(s)P_{p}(s)$ is employed. Then, substitute $P$ with $P_{p}$ in (29) and by letting $\epsilon_{i}\to 0$ solve (III-C) to get stabilizing $K_{p}(s)\in\mathbb{R}(s)$, virtual loop’s control system over the scalar approximation $P_{P}(s)$. Then the scalar complementary sensitivity function for this modified virtual loop is given by
$T{\color[rgb]{0,0,0}{}_{p}}(s)\triangleq\Bigr{(}P_{p}(s)K_{p}(s)\Bigl{)}/\Bigr{(}1+P_{p}(s)K_{p}(s)\Bigl{)}$ (36)
Theorem 4.
Consider $P(s)$ as defined in (11)-(12) with $\text{rank}_{\mathbb{R}(s)}(P)=n_{u}<n_{y}$ and $Z_{R}\neq\emptyset$. Then, the approximate inverse is
$\Xi_{a}^{\ell}(s)=T_{p}(s)P^{\dagger}(s).$ (37)
and the corresponding control input is
$u_{inv}(t)=\mathscr{L}^{-1}\{T{\color[rgb]{0,0,0}{}_{p}}(s)P^{\dagger}(s)Y(s)\}.$ (38)
Assuming $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$, then $u_{inv}(t)$ satisfies
$\|e(t)\|\leq\alpha\|e(0)\|e^{-\beta t},\forall t>0,$ (39)
for some $\alpha,\beta>0$. while ensuring $\|u_{inv}(t)\|_{\infty}<\infty$.
Proof.
Note that $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)\implies U_{R1}\hat{U}_{R1}{Y}(s)={Y}(s)\implies P(s)P^{\dagger}(s)Y(s)=Y(s)$ and $P^{\dagger}(s)$, as in Corollary 2 solves (11) but with an unstable $u_{inv}(t)$ since $Z_{R}\neq\emptyset$. To get a stable $u_{inv}(t)$, we can substitute $T_{p}(s)P^{\dagger}(s)Y(s)$ into (33) as
$E(s)=Y(s)-P(s)T_{p}(s)P^{\dagger}(s)Y(s)$
and since $T_{p}(s)$ is scalar, using $P(s)P^{\dagger}(s)Y(s)=Y(s)$, and $S_{p}(s)\triangleq I-T_{p}(s)$, we get $E(s)=Y(s)S_{p}(s)$.
Assume $p$ is an unstable pole of $Y(s)$. Then since $K_{p}$ in (36) is obtained by (III-C), which gives $\begin{bmatrix}I&K_{p}\\ -P_{p}&I\end{bmatrix}^{-1}\in\mathcal{R}\mathcal{H}_{\infty}$, $P(s)$ and $P_{p}(s)$ shares the same poles, we have $S_{p}(s)\in\mathcal{RH}_{\infty}$ and $S_{p}(p)=0$ and then following the proof of Theorem 3 yields $\|e(t)\|\leq\alpha\|e(0)\|e^{-\beta t},\quad\forall t>0.$ Finally, Lemma 1.(1) applies on $T_{p}(s)$ so that $T_{p}(p)=0$ yielding $\|u_{inv}(t)\|_{\infty}<\infty$. This completes proof. ∎
Remark 4.
Although operators such as $S_{i}(s)K(s)$, $S_{i}(s)$, and $T_{i}(s)$ are commonly used to analyze standard closed-loop responses, here we repurpose $S_{i}(s)K(s)$ (in Theorem 3) or $T_{p}(s)$ (in Theorem 4) as a direct control action - rather than using $K(s)$ solely like a conventional loop, which departs from typical closed-loop treatments. Consequently, while alternative methods can also be used to design the $K(s)$ and therefore $S_{i}(s)K(s)$ and $T_{p}(s)$ the $\mathcal{H}_{\infty}$ framework provides valuable analytical properties. In particular, it yields an “optimal” approximate inverse within our setting.
Remark 5.
Theorem 3 and Theorem 4 provide stable approximations if the solution in (III-A) is unstable. The control input $u_{inv}(t)$ in both theorems is not unique, allowing design flexibility via (29). In Theorem 3, $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$ always holds as $P(s)$ is square and full rank, while relaxing this condition, $Y(s)\notin\operatorname*{Im}_{\mathbb{R}(s)}(P)$, in Theorem 4 introduces a non-zero $\varphi$ in (39), as shown in (9).
All results assume a nominal system $P(s)$, but real-world complexities such as modeling errors and disturbances can lead to large errors or instabilities. These implications, along with structures over $\varphi$, will be discussed in the next section.
III-D Robustness
Except for a few studies [16], to the best of our knowledge, investigations regarding stable inversion are typically conducted assuming that a complete understanding of the model is available which is often not feasible. In this section, we extend the results of Section III-C by further relaxing the constraints set forth in Theorem 1. Specifically, Theorem 5 addresses how the condition $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$ can be relaxed under uncertainties, while Corollaries 4 and 5 explore the full range of possible relaxation scenarios. For the desired output, (11)-(12) can be rewritten as
$P(s)U(s)=Y_{d}(s)$ (40)
where $Y_{d}(s)$ denotes the desired output. Now, to align the tracking problem in (40) to the results obtained in Sections III-B and III-C the following Assumption is useful.
Assumption 2.
The desired trajectory $y_{d}(t)\triangleq\mathscr{L}^{-1}\{Y_{d}(s)\}$ and $P(s)$ satisfies:
  • (1)
    ${\color[rgb]{0,0,0}Y_{d}(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)\cap\mathcal{RH}_{\infty}}$,
  • (2)
    If $z$ is a RHP zero of $P(s)$, i.e., $z\in\mathbb{C}$ s.t. $\Re(z)>0$ and $y^{*}_{z}P(z)=0$, then $Y_{d}(z)=0$,
  • (3)
    $P(s)$ is stable or stabilized.
Now we can consider uncertainties over $P(s)$ in (40). Let the perturbed plant, $P_{\Pi}(s)$, be a member of all possible plants
$P_{\Pi}(s)\in\Pi\triangleq\left\{({I}+W_{1}\Delta_{u}{W}_{2}){P}\right\}$ (41)
where $W_{1},W_{2}$ are TFMs that characterize the spatial and frequency structure of the uncertainty, $\Delta_{u}$ denotes any unknown unstructured function with $\|\Delta_{u}\|_{\infty}<1$ [27, chapter 8.1]. Moreover, the perturbed (real) system’s zeros are
$\text{Z}_{P_{\Pi}}$$\triangleq\{z\in\mathbb{C}:P_{\Pi}(z)u_{z}=0y_{z}\}$ (42)
$\text{Z}_{R_{\Pi}}$$\triangleq\{z\in\text{Z}_{P_{\Pi}}:\Re(z)>0\}\subseteq\text{Z}_{P_{\Pi}}.$ (43)
Here ${Z}_{P_{\Pi}}$ might be different than $Z_{P}$ which means $P_{\Pi}$ in (41) considers uncertain RHP zeros for $P(s)$. Then, the perturbed output for a given input as
$y_{\Pi}(t)=\mathscr{L}^{-1}\{P_{\Pi}(s)U(s)\}=\mathscr{L}^{-1}\{Y_{\Pi}(s)\}.$ (44)
Note that, the inner product for any vector-valued functions $F,G\in\mathcal{L}_{2}(j\mathbb{R})\cap\mathbb{R}(s)^{n_{y}}$ is defined as:
$\langle F,G\rangle\triangleq\frac{1}{2\pi}\int_{-\infty}^{\infty}\sum_{k=1}^{n_{y}}F_{k}^{\ast}(j\omega)G_{k}(j\omega)d\omega$ (45)
Consider $L$ given in (2), as $P(s)\in\mathcal{L}_{2}(j\mathbb{R})\cap\mathbb{R}(s)^{n_{y}\times n_{u}}$, $L\in\mathcal{L}_{2}(j\mathbb{R})$. Given that the columns in $L$ are not necessarily orthogonal, we can rectify this by applying the Gram-Schmidt process to the columns of $L$. This process can be expressed as:
$v_{i}=p_{i}-\sum_{j=1}^{i-1}\frac{\langle p_{i},q_{j}\rangle}{\langle q_{j},q_{j}\rangle}q_{j},q_{i}=\frac{v_{i}}{\sqrt{\langle v_{i},v_{i}\rangle}},i=1,\ldots,r.$ (46)
Here, the $v_{i}$ vectors are orthogonal, and the $q_{i}$ vectors are orthonormal. Based on this, the orthonormal set is defined as ${Q}=[q_{1}\ \dots\ q_{r}]\in\mathcal{L}_{2}(j\mathbb{R})\cap\mathbb{R}(s)^{n_{y}\times r}$ which spans the same subspace as $L$ in (2). With these conditions the transformation from $L$ to $Q$ always exists.
Definition 1.
Let $\operatorname*{Im}_{\mathbb{R}(s)}(P)$ is spanned by an orthonormal basis $\{q_{k}\}_{k=1}^{r}\subset\mathcal{L}_{2}(j\mathbb{R})$. For any $F\in\mathcal{L}_{2}(j\mathbb{R})$, the projection onto $\operatorname*{Im}_{\mathbb{R}(s)}(P)$ is given by:
$\operatorname*{proj}_{\operatorname*{Im}{(P)}}[$ $F]=\sum_{i=1}^{r}\langle F,q_{i}\rangle q_{i},$ (47)
where $f$ can be decomposed as:
$F=\operatorname*{proj}_{\operatorname*{Im}(P)}[F]+\operatorname*{res}[F],$ (48)
with $\langle\operatorname*{res}[F],q_{k}\rangle=0,\forall k\in\{1,\dots,r\}.$ (49)
Then, the overall feedback strategy combining two feedforward actions to deal with uncertainties is given by Fig. 1. Here, $u_{ff}(t)$ is feedforward control input, $u_{fb}(t)$ is feedback control input, and $u_{c}(t)$ is combined (effective) control input $u_{c}(t)$, $y_{\Pi}(t)$ denotes the real output under $u_{c}(t)$, and we have $y_{\Delta}(t)\triangleq y_{\Pi}(t)-y(t)$.
Based on the conditions, exact inversion (Fig. 1(c)(i)) is valid for all system classes, including non-minimum phase, square, non-square, and singular systems, though it requires an output redefinition in the feedback path, $\bar{y}_{\Delta}(t)=({z^{+}}/{z_{d}^{+}})\ast y_{\Delta}(t)$ where $z^{+}(s)$ given by (22) and the denominator is
$z^{+}_{d}(s)=\prod_{z\in Z_{R}}(s+z)^{m_{z}}\prod_{z\in Z_{C}}(s+z)^{m_{z}}(s+{z}^{*})^{m_{{z}^{*}}}.$ (50)
Approximate inversion is applicable to non-minimum phase, full-rank systems; for square systems, the loop corresponds to Fig. 1(c)(ii), and for non-square systems, it corresponds to Fig. 1(c)(iii). Then, over the nominal system $P(s)$, the feedforward control signal $u_{ff}(t)$ is obtained by solving (III-A) (Theorem 1), (31) (Theorem 3), or (37) (Theorem 4), subject to (40). Similarly, $u_{fb}(t)$ and $u_{ff}(t)$ are obtained by following the same procedures. The next question, whether the proposed loop in Fig.1 can compensate the error caused by uncertainty, is revealed by the upcoming theorem holding P1-P2 and Assumption 2.
[…]
[width=]blckdiag.pdf
Figure 1: Block diagram for closing the loop for stable inversion.
Theorem 5.
Let the control inputs in Fig. 1 be defined as $u_{ff}(t)=\Xi_{i}^{\ell}(y_{d}(t))$ and $u_{fb}(t)=\Xi_{i}^{\ell}(\bar{y}_{\Delta}(t))$ and designed over $P$ in (11) with $Z_{R}\neq\emptyset$. Suppose the actual system $P_{\Pi}$ be defined as in (41) with $P_{\Pi}\neq P$. For the combined control law $u_{c}(t)=u_{ff}(t)-u_{fb}(t)$, the resulting tracking error satisfies
$\bigl{(}y_{d}(t)-{y}_{\Pi}(t)\bigr{)}\to\mathscr{L}^{-1}\{\text{res}[Y_{\Pi}]\}$ (51)
Proof.
For brevity, let us denote $W_{1}\Delta_{u}{W}_{2}$ simply by $\Delta_{u}$. Although $\operatorname*{rank}_{\mathbb{R}(s)}({P}_{\Pi})\leq\operatorname*{rank}_{\mathbb{R}(s)}({P})$, it may still occur that $\text{Im}_{\mathbb{R}(s)}(P_{\Pi})\not\subseteq\text{Im}_{\mathbb{R}(s)}(P)$ implying $\text{res}[Y_{\Pi}]=\text{res}[Y_{\Delta}]\neq 0$. Following [37, pp. 283–294], we adjust the Laplace transform integral to circumvent algebraic loops. Specifically, we define:
[…]
where $\epsilon>0$ can be taken arbitrarily small which ensures that the Laplace transform is taken in a way that sidesteps any instantaneous feedback algebraic loop. Consider now the closed-loop configuration depicted in Figure 1, as
$U_{fb}(s)$ $=\Xi_{i}^{\ell}(s)\bar{Y}_{\Delta}(s)=\frac{z^{+}(s)}{z_{d}^{+}(s)}\Xi_{i}^{\ell}(s)Y_{\Delta}(s)$ (53)
$=\frac{z^{+}(s)}{z_{d}^{+}(s)}\Xi_{i}^{\ell}(s)\bigr{(}\operatorname*{proj}_{\operatorname*{Im}(P)}[Y_{\Delta}]+\operatorname*{res}[Y_{\Delta}]\bigl{)}$ (54)
where $\langle\operatorname*{res}[Y_{\Delta}],q_{k}\rangle=0$ implying $\Xi_{i}^{\ell}(s)\operatorname*{res}[Y_{\Delta}]=0$. Thus, the feedback only responds to the part of $Y_{\Delta}$ in $\operatorname*{Im}_{\mathbb{R}(s)}(P)$. Consequently,
$PU_{fb}=\frac{z^{+}(s)}{z_{d}^{+}(s)}\operatorname*{proj}_{\operatorname*{Im}(P)}[Y_{\Delta}]$ (55)
Next, we express the integral equation over $Y_{\Delta}$:
$(I+\Delta_{u})P(s)U_{c}(s)-P(s)U_{c}(s)=\Delta_{u}P(s)U_{c}(s)$ (56)
$=Y_{\Delta}(s)=\underbrace{\Delta_{u}P(s)U_{ff}(s)}_{\begin{subarray}{c}\in\mathcal{RH}_{\infty}\text{ and}\\ \text{ invariant under }u_{fb}\end{subarray}}-\Delta_{u}P(s)U_{fb}(s).$ (57)
then using (53) yields
$\Delta_{u}P(s)U_{ff}(s)-\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\ell}(s)Y_{\Delta}(s)=Y_{\Delta}(s)$
$\Delta_{u}P(s)U_{ff}(s)=\Bigl{(}I+\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\ell}(s)\Bigr{)}Y_{\Delta}(s)$
Since $P(I-V_{R1}\hat{V}_{R1})\bar{\kappa}=0$, $\max_{\omega}\bar{\sigma}\Bigl{(}P\hat{V}_{R}\begin{bmatrix}\Lambda^{-1}(s)&\textbf{0}\\ \textbf{0}&\textbf{0}\end{bmatrix}\hat{U}_{R}\Bigr{)}=1$, $\biggl{\|}\frac{z^{+}(s)}{z_{d}^{+}(s)}\biggr{\|}_{\infty}=1$, and $\|\Delta_{u}\|_{\infty}<1$, it follows that:
$\biggl{\|}\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\ell}(s))\biggr{\|}_{\infty}=\alpha_{b}<1$
which ensures convergence of the following Neumann series.
$\sum_{k=0}^{\infty}\biggl{(}\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\ell}(s)\biggr{)}^{k}$ (58)
Thus, $\Bigl{(}I+\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\ell}(s)\Bigr{)}^{-1}$ exists and stable with $\|Y_{\Delta}\|_{\infty}\leq\frac{\|\Delta_{u}P(s)U_{ff}(s)\|_{\infty}}{1-\alpha_{b}}$. Then, the tracking error is
$Y_{d}(s)-Y_{\Pi}(s)$ $=Y_{d}(s)-Y_{\Delta}(s)-\underbrace{P(s)U_{ff}(s)}_{\triangleq Y_{d}(s)}+P(s)U_{fb}(s)$
$=-Y_{\Delta}+P(s)U_{fb}(s)$
$=-\operatorname*{proj}_{\operatorname*{Im}(P)}[Y_{\Delta}]-\operatorname*{res}[Y_{\Delta}]+\frac{z^{+}(s)}{z_{d}^{+}(s)}\operatorname*{proj}_{\operatorname*{Im}(P)}[Y_{\Delta}]$
Finally, $\mathrm{res}[Y_{\Pi}]=\mathrm{res}[Y_{\Delta}]$, and since
$\lim_{s\to 0}\biggl{[}s\biggl{(}I-\frac{z^{+}(s)}{z_{d}^{+}(s)}\biggr{)}\operatorname*{proj}_{\operatorname*{Im}(P)}[Y_{\Delta}]\biggr{]}=0,$
standard final value arguments imply that as $t\to\infty$, the components in $\operatorname*{Im}_{\mathbb{R}(s)}(P)$ converges to zero. Hence,
$\lim_{t\to\infty}\bigl{(}y_{d}(t)-y_{\Pi}(t)\bigr{)}=\mathcal{L}^{-1}\bigl{\{}\mathrm{res}[Y_{\Pi}]\bigr{\}}.$
This shows that the tracking error converges to the inverse Laplace transform of the non-cancellable yet stable residual term, thus completing the proof.
Corollary 4.
Assume $P(s)$ is square, full rank implying $\operatorname*{res}[Y_{\Pi}]=0$, and $Z_{R}\neq\emptyset$, where the control inputs be defined as $U_{ff}(s)=S_{i}(s)K(s)Y_{d}(s)$ and $U_{fb}(s)=S_{i}(s)K(s)\bar{Y}_{\Delta}(s)$. Then (51) becomes $\bigl{(}y_{d}(t)-{y}_{\Pi}(t)\bigr{)}\to 0$.
Corollary 5.
For the depicted block diagram in Fig. 1 with following conditions, (51) can also be re-written:
[…]
$\implies\bigl{(}y_{d}(t)-y_{\Pi}(t)\bigr{)}=0$
[…]
$\implies\bigl{(}y_{d}(t)-y_{\Pi}(t)\bigr{)}\to 0$
[…]
$\implies\bigl{(}y_{d}(t)-y_{\Pi}(t)\bigr{)}=\mathscr{L}^{-1}\{\text{res}[Y_{\Pi}]\}\ \forall t>0$
Here, the term $\mathscr{L}^{-1}\{\text{res}[.]\}$ corresponds directly to $\varphi$ as described in (9), representing the contribution of inevitable errors. On the other hand, from an implementation perspective, the compactness of the algebraic structures allows for explicitly stating the existence and boundedness conditions and facilitates straightforward solutions of stable inversion. However, when utilizing $\mathcal{H}_{\infty}$-based approximate solutions, we encounter complex, high but finite-order algebraic structures. These structures should be reduced to an acceptable order through balanced model reduction techniques.
IV Numerical Examples
In this section, we present several numerical examples to illustrate the practical application and effectiveness of our proposed approach for model inversion. These examples cover various scenarios demonstrating the broad applicability and flexibility of the proposed method. Lastly, these examples aim to provide an intuitive understanding of the theoretical concepts previously discussed.
Example 1.
Consider the following state-space representation.
$\dot{x}(t)$ $=\begin{bmatrix}-3.565&-0.1919&0.09499\\ 4.418&-1.44&-0.0847\\ -0.3121&0.157&-0.005135\end{bmatrix}x(t)$ (59)
$\quad+\begin{bmatrix}-1.555&0.4156\\ 1.173&0.4576\\ -0.05738&0.1017\end{bmatrix}u(t)$
$y(t)$ $=\begin{bmatrix}-0.2573&0.4612&-1.024\\ 0.2573&-0.4612&1.024\end{bmatrix}x(t)$
where the two-input-two-output system is minimal, and has an invariant zero at $s=1$ (so it is a non-minimum phase). Now define $P(s)\triangleq C(sI-A)^{-1}B$. Then, $\operatorname*{rank}_{\mathbb{R}(s)}(P(s))=1<n_{u}=n_{y}$, leading to a so-called singular MIMO that fits Theorem 1. Next, consider the desired output $Y_{d}(s)\triangleq\begin{bmatrix}{1}/{s}&-{1}/{s}\end{bmatrix}^{T}$. In accordance with Assumption 2, we modify $Y_{d}(s)$ to be $Y_{d}(s)=\frac{s-1}{0.5s+1}\begin{bmatrix}\frac{1}{s}&-\frac{1}{s}\end{bmatrix}^{T}$. This makes $Y_{d}(s)$ exactly trackable such that $\operatorname*{rank}_{\mathbb{R}(s)}\left[P(s)\right]=\operatorname*{rank}_{\mathbb{R}(s)}\left[P(s):Y_{d}(s)\right]=1$. Then the resulting control input $u(t)$ in (59) is
$U(s)=\begin{bmatrix}\frac{2s+6}{s}-\frac{100}{100s+1}\\ 1\end{bmatrix}$ (60)
then to eliminate the Dirac-delta functions resulting from inverse Laplace transformation, multiplying (60) with a suitable low-pass filter and taking the inverse Laplace transformation yields
$u(t)=\mathscr{L}^{-1}\Bigl{\{}U(s)\frac{1}{0.01s+1}\Bigr{\}}=\begin{bmatrix}195e^{-100t}-e^{-0.01t}+6\\ 100e^{-100t}\end{bmatrix}$
The tracking result is given by Fig. 2. It is shown that both outputs track the reference signals which validates the theory where $y_{d}(t)=y(t)$.
[…]
[width=0.9]ex1_outs
Figure 2: The output tracking of Example 1 as Theorem 1.
Example 2.
As a comparison to model matching problem in [11], consider the following $2\times 2$, full rank and minimal system that has an invariant zero at $s=1$.
$P(s)=\begin{bmatrix}\frac{(1-s)}{(s+1)^{2}}&\frac{0.3}{s+0.5}\\ \frac{-(1-s)}{(s+1)^{2}(s+2)}&\frac{2}{s+3}\end{bmatrix}$ (61)
In this case, $\operatorname*{rank}_{\mathbb{R}(s)}(P(s))=2=n_{u}=n_{y}$, which implies that all possible output combinations can be tracked. Given Theorem 3, the modifications introduced by Assumption 2.(2) are unnecessary. Therefore, let’s consider $Y_{d}(s)=\begin{bmatrix}{1}/{s}&{1}/{s}\end{bmatrix}^{T}$, which results in $\operatorname*{rank}_{\mathbb{R}(s)}\left[P(s)\right]=\operatorname*{rank}_{\mathbb{R}(s)}\left[P(s):Y_{d}(s)\right]=2$. We can solve equation (III-C) for (61) with $\gamma=3.6$. As $S_{i}(s)K(s)$ is strictly proper, this is all in line with Theorem 3. From the solution, it can be observed that both Theorem 2 and Lemma 1.(1) are satisfied. Consequently, the resulting $u(t)\triangleq\mathscr{L}^{-1}\{U(s)\}$ in equation (31) is depicted in Fig. 3 and resulting $S_{i}(s)K(s)$ and $T_{i}(s)$ are given by Fig. 4. Compared with the results obtained using the Inner-Outer Factorization method (refer to [11, Fig.8 - black lines]), the proposed method considerably enhances tracking performance (almost at par with the feedback case) in the feedforward scenario.
[…]
[width=0.9]ex2_outsandinps.pdf
Figure 3: The output tracking of Example 2 and corresponding inputs as Theorem 3.
[…]
[width=0.9]ex2_SKT.pdf
Figure 4: The singular value plots of $S_{i}(s)K(s)$ and $T_{i}(s)$.
Example 3.
Now, consider the following $n_{y}=2,n_{u}=1$ system which is an example of an underactuated system.
$P(s)=\begin{bmatrix}\frac{s^{2}-5s-50}{s^{2}+3s+2}\\ \frac{s-10}{s^{2}+3s+2}\end{bmatrix}$ (62)
where the system has an invariant zero at $s=10$, thus it is also a non-minimum phase system. Now since choosing $Y_{d}\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$, is much straightforward, we set $Y_{d}$ as
$Y_{d}(s)=\begin{bmatrix}\frac{2}{s^{2}+1}+\frac{8}{s}\\ \frac{1}{s}\end{bmatrix}\notin\operatorname*{Im}_{\mathbb{R}(s)}(P)\text{ s.t.}\operatorname*{res}(Y_{d})\neq 0$ (63)
Then, let’s choose the scalar $P_{p}(s)=\frac{s^{2}-5s-50}{s^{2}+3s+2}$ which also has an invariant zero at $s=10$ and thus no need to augmentation of $z^{+}$, and then solving (III-C) yields a stable $T_{p}(s)$ with $\gamma=2.32$. Also, based on (62), we have
$P^{\dagger}(s)=\begin{bmatrix}\frac{s^{3}+8s^{2}+17s+10}{s^{3}-74s-260}&\frac{s^{2}+3s+2}{s^{3}-74s-260}\end{bmatrix}$ (64)
Then using Theorem 4, we have a strictly proper controller resulting in $T_{p}(s)P^{\dagger}(s)$. The reference tracking performance of the form in Theorem 4 shown by Fig. 5 and the design of $T_{p}(s)$ and the boundedness of $u(t)$ are given by Fig. 6. As it can be seen from Fig. 5, the errors are fully in harmony with Definition 1.
[…]
[width=0.9]ex3_outs.pdf
Figure 5: The output tracking of Example 3 as Theorem 4.
[…]
[width=0.9]ex3_inpandT.pdf
Figure 6: The control input and singular values of $T(s)$ of Example 3.
Example 4.
We again consider the same nominal system as in (61) whose invariant zeros are located at $s_{1,2,3}=-10,-0.86,1$ and poles at $s_{1,2,3,4}=-1,-1,-0.5,-3$. Then keep $S_{i}(s)K(s)$ of Example 2 as a controller for $\gamma=3.6$ in Fig. 4. Then, $S_{i}(s)K(s)$ is used both for feedforward and feedback actions in Fig. 1. Then, we start perturbating the system resulting as
${P}_{\Pi}(s)=\begin{bmatrix}\frac{(0.2-s)}{s^{2}+20s+100}&\frac{s^{2}+1.8s-0.4}{s^{2}+8s+16}\\ \frac{0.3}{s+0.1}&\frac{2}{s+2.1}\end{bmatrix}$ (65)
whose invariant zeros are located at $s_{1,2,3,4,5}=-20.4,-6.4,-2.7,-1.2,0.2$ and poles at $s_{1,2,3,4,5,6}=-10,-10,-4,-4,-0.1,-2.1$.
Example 5.
Consider the following state-space representation.
$\dot{x}(t)$ $=\begin{bmatrix}-0.8808&1.876&-1.011&0.8287\\ 0.1876&-0.5055&-0.0869&-0.8554\\ 0&0&-0.9163&0\\ 0&0&0&-0.5108\end{bmatrix}x(t)$ (66)
$\quad+\begin{bmatrix}2.196&-1.785\\ -0.1832&1.813\\ 1.527&0\\ 0&1.277\end{bmatrix}u(t)=Ax(t)+Bu(t)$
$y(t)$ $=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\end{bmatrix}x(t)=Cx(t)$
Equation (66) discrete form with a sample time of $1$ seconds is given by [15], geometric approaches in multivariable control, which is a benchmark in the literature. $C(sI-A)^{-1}B$ is a full-rank, minimal, and square system implying all possible desired output can be achieved whether bounded or unbounded control inputs. Moreover, $Z_{P}=\{-0.2149,0.0925\}$, $Z_{R}=\{0.0925\}$ indicating the system has a dominant RHP zero making the system non-minimum phase. The reference signal is
$y_{d}(t)=\begin{bmatrix}y_{d1}(t)\\ y_{d2}(t)\end{bmatrix}=\begin{cases}[\frac{t}{20}\ 0]^{T}&\text{ for }0\leq t<20\\ [1\ 0]^{T}&\text{ for }t\geq 20\end{cases}$ (67)
$Y_{d}(s)=\begin{bmatrix}(1-e^{-20s})/(20s^{2})\\ 0\end{bmatrix}\text{ s.t. }Y_{d}(0.0925)\neq\begin{bmatrix}0\\ 0\end{bmatrix}$ (68)
Given conditions show that Theorem 3 applies for the feedforward part in Fig.1 with setting $u_{fb}(t)=0$. We can solve equation (III-C) for (66) with a finite $\gamma$ where resulting $S_{i}(s)K(s)$ and $T_{i}(s)$ functions, given by Fig. 7, are strictly proper. This is all in line with Theorem 3.
[…]
[width=0.9]ex6_SKT.pdf
Figure 7: The singular value plots of $S_{i}(s)K(s)$ and $T_{i}(s)$ for Example 5.
After that, the controller $U^{a}(s)$ is discretized by Bilinear transform with $1$ seconds of sample time in order to align the results to the ones in [15, Section 5]. The discrete desired outputs, $y_{d}(t)$, the outputs, $y_{1}(t),y_{2}(t)$-proposed, with respect to the implemented controller $U^{a}(s)$ in this proposed approach, and the outputs, $y_{1}(t),y_{2}(t)$-geometric, under geometric control are given in Fig. 8.
[…]
[width=]ex6_outs.pdf
Figure 8: The output tracking comparison of approximated control in Theorem 3 to geometric approaches in [15].
Finally, the corresponding inputs are given in Fig. 9.
[…]
[width=0.9]ex6_ins.pdf .
Figure 9: The control inputs produced by $U^{a}(s)$
From the results, it is evident that the 20th-degree $U^{a}(s)$ obtained with Theorem 3 and discretized by using Tustin transformation provides almost the same tracking performance as the geometric approach. The most important detail here is that the computation of $U^{a}(s)$ does not require a 100-second preview used in geometric control which opposes causality. Additionally, the structure in (III-C) offers design flexibilities for practitioners to adjust performance, usage, and size for implementation.
V Conclusion
In this paper, we present a comprehensive approach to stable model inversion aiming to establish necessary and sufficient conditions for stable model inversion using algebraic control.
The proposed approach in this paper applies to various system types, including square, non-square, rank-deficient, and non-minimum phase systems. It is validated by numerical examples is discussed. Our method’s robustness against uncertainties encountered in practical applications expands the model inversion framework and surpasses limitations in existing methods.